OpenStudy (anonymous):

Solve by elimination 5x+6y=1 10x+12y=2

OpenStudy (anonymous):

both the equations are the same.

OpenStudy (anonymous):

What do you mean??

OpenStudy (anonymous):

Theyre the same line,

OpenStudy (diyadiya):

5x+6y=1 multiplied by 2 =10x+12y=2

OpenStudy (anonymous):

so is there no solution or many solutions?

OpenStudy (anonymous):

infinitely many solutions ... multiply by 2 the first : 10x+12=2 now you can subtract one from another and get 0=0 which is true

OpenStudy (anonymous):

aaahhhh I see how you did that now thanks :D